# A Configuration-First Framework for Reproducible, Low-Code Localization

### v2

@@ Abstract
As machine learning (ML) becomes the backbone of critical applications—ranging from logistics to healthcare it becomes even more important to keep the results credible and comparable. Everyday workflows should make rigorous experiment specification and exact repeatability the default, without blocking advanced experimentation. However, in practice, researchers face a three-way gap that could be filled by a framework that offers (i) low coding effort for end-to-end studies, (ii) reproducibility by default including versioned code/data/configurations, controlled randomness, isolated runs, recorded artifacts, and (iii) built-in extensibility so new models, metrics, and stages can be added with minimal integration effort. Existing tools rarely deliver all three for machine learning in general and localization workflows in particular. In this paper we introduce LOCALIZE, a low-code, configuration-first framework for radio localization in which 1) experiments are declared in human-readable configuration, 2) a workflow orchestrator runs standardized pipelines from data preparation to reporting, and 3) all artifacts, such as datasets, models, metrics, and reports are versioned. The pre-configured, versioned datasets reduce initial setup and boilerplate speeding up model development and evaluation. The design, with clear extension points, let experts add components without reworking the infrastructure. In a qualitative comparison and a head-to-head study against a plain Jupyter notebook baseline, we show that the framework reduces authoring effort while maintaining comparable runtime and memory behavior. Furthermore, using a Bluetooth low-energy dataset we show that scaling across training data ($1\times$ to $10\times$) keeps orchestration overheads bounded as data grows. Overall, the framework makes reproducible machine-learning-based localization experimentation practical, accessible, and extensible.
@@ I Introduction
Background and motivation. The engineering of modern data-driven systems has undergone a paradigm shift. As machine learning (ML) becomes the backbone of critical applications—ranging from logistics to healthcare—the software engineering (SE) community faces a growing challenge: the reproducibility crisis in research software [Pineau2021, Heil2021]. While ML provides powerful mappings from raw data to actionable insights, the development process remains largely ad-hoc, characterized by fragile codebases where small changes in global state or environment lead to inconsistent results [bouthillier_accounting_2021]. This issue is particularly acute in radio-based localization powering the majority of location-based applications such as navigation, logistics, social applications, entertainment, and healthcare [huang2018location]. As the radio-based localization research and engineering is transitioning from rigid analytical models to ML powered ones [trevlakis_localization_2023], the research and development tools and pipelines that must evolve to support this evolution.
@@
Gaps in current practice. Moving from analytical to machine learning approaches has shown promise in many settings, but raises practical challenges for everyday research. The complexity and heterogeneity of ML pipelines raise the barrier to entry as shown in a recent exploration of AI system development [dolata2023making]. The study highlights that the stochastic nature of ML introduces a unique cognitive load on developers that traditional SE practices do not adequately address. GUI/low-code platforms reduce coding effort, but can constrain advanced experimentation and fine-grained control [weka_book, orange_pdf]. Conversely, code-centric workflows provide maximum flexibility but, absent dedicated tooling, place the burden of reproducibility, environment control, and consistent evaluation on the researcher [Pineau2021, Heil2021, henderson_deep_2018, bouthillier_accounting_2021, gundersen2021do]. In many localization studies, the primary novelty is a new model or training method evaluated on public datasets [trevlakis_localization_2023]. Onboarding data and recreating baselines becomes setup overhead rather than the research focus. The broader ML literature shows that small changes to preprocessing, hyperparameters, software versions, or random seeds can alter conclusions and hinder fair comparison [Pineau2021, Heil2021]. Beyond statistical variance, exact repeatability is fragile: hidden global state, non-deterministic operations, data-order effects, and environment or hardware differences can change outputs even when seeds and settings are fixed [bouthillier_accounting_2021, henderson_deep_2018, pham_problems_nodate, 10855627]. Such issues are easy to miss and hard to diagnose.
@@
This challenge is not unique to radio localization but represents a critical bottleneck in many scientific domains now adopting machine learning [Heil2021]. Fields such as computational biology and materials science face a similar gap where the transition from traditional models to complex, data-driven pipelines strains existing research practices [Heil2021]. The lack of standardized, reproducible, and extensible workflows leads to duplicated effort, fragile results, and a high barrier to entry for domain experts who are not also software engineers [wilson2017good]. Therefore, the framework presented in this paper addresses a fundamental software engineering problem: how to engineer research-specific development environments that enforce scientific rigor by default without sacrificing the agility needed for rapid experimentation.
@@
Proposed approach and contributions. To enable the development of high-quality generic models, the wireless localization research community could benefit from a well engineered framework enabling: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) flexibility that is built in, so advanced experimentation can be conducted without fighting the tool-chain. While many platforms address ease of use or extensibility, few make reproducibility the default alongside both. This missing combination is the focus of our work and follows the principles within engineering within boundaries [10883021]. This paper introduces LOCALIZE  [^1] https://github.com/sensorlab/localize, a low-code, configuration-first framework that makes ML-based radio localization practical, accessible, and reproducible by reducing initial effort, enabling novice and expert use, and supporting expert extensibility while promoting sound experimental practices. The contributions of this paper presents are as follows:
@@
  • We introduce LOCALIZE, a low-code, configuration-first framework that bridges ease of use with scientific rigor by making reproducibility its default operating mode, integrating version control, execution isolation, and transparent artifact tracking.
@@
  • We design LOCALIZE to enforce consistent, comparable evaluation by applying a standardized set of metrics and reporting procedures across all methods and datasets, eliminating glue code and improving the credibility of results.
@@
In the remainder of this paper, we review related work in Section II, outline the problem and requirements in Section III, describe the proposed framework in Section IV, present a reference implementation in Section V, and evaluate it against established platforms in Section VII. Finally, Section IX concludes the paper.
@@ II-A The Challenge of Reproducibility in Machine Learning
A consistent lesson from the literature is that ML results can be sensitive to seemingly minor choices, so reproducibility requires attention to both variability and exact repeatability. Bouthillier et al. [bouthillier_accounting_2021] quantify sources of variability in benchmarks and show that initialization, data sampling and order, and hyperparameter settings all affect outcomes. They recommend practices such as repeated runs and clearer reporting to separate algorithmic gains from noise. Building on this empirical picture, Henderson et al. [henderson_deep_2018] demonstrate that non-determinism in benchmark environments and implementation details can materially change results, arguing for stricter methodology and documentation. Furthermore, [10855627] recently formalized the concept of ’reproducibility smells’ defined as patterns in the development lifecycle, such as hidden global states and lack of automated artifact tracking, that lead to fragile and non-deterministic results. Collectively, these studies establish that achieving computational reproducibility requires meticulous control over the entire experimental pipeline.
In response, subsequent work outlines reporting and release requirements. Heil et al. [Heil2021] introduce tiered levels of computational reproducibility (bronze/silver/gold) with concrete requirements covering code, data, captured environments, and automation. At the community level, Pineau et al. [Pineau2021] describe the NeurIPS reproducibility program, checklists, a code policy, and a community challenge, which raised expectations for transparency across submissions.
Despite these advances, practical support in general ML platforms remains uneven. Gundersen et al. [gundersen2021do] report that reproducing published results often requires non-trivial user effort. This suggests a clear need for research workflows that embed controls, documentation, and repeatable execution as default, integral features rather than optional, post-hoc additions.
@@ II-B General-Purpose Experimentation Platforms
Several general-purpose platforms aim to simplify ML experimentation while only partly alleviating the cognitive load of AI development [dolata2023making]. They typically fail to resolve the tension between usability, reproducibility, and extensibility, forcing researchers to make compromises. For example, graphical tools like WEKA [weka_book] and visual programming environments like Orange3 [orange_pdf] excel at lowering the barrier to entry in machine learning research in different domains. They allow for rapid prototyping without extensive coding. However, they externalize the burden of reproducibility, as rigorously capturing an end-to-end experiment, including precise library versions and sources of non-determinism, requires manual effort outside the core workflow.
Declarative, code-centric platforms like Ludwig [ludwig_pdf] advance this goal by using configuration files to define deep learning experiments. This approach streamlines workflows and supports extensibility. While powerful for neural networks, it is not primarily designed for the classical machine learning pipelines common in localization studies and, similar to graphical tools, leaves environment and data versioning as an external concern.
These platforms assist with parts of the problem—addressing low-code use or standardizing certain workflows, but they do not jointly provide a solution that is reproducible by default, accessible via a low-code interface, and flexible enough for advanced experimentation.
@@ II-C Other Related Efforts
Specialized tools also contribute valuable insights, typically focusing on specific domains or aspects of experimental management. Some efforts focus on automating workflows for a specific domain, such as CSI fingerprinting localization [Tong2021] or multivariate time-series forecasting [ye2024mvts], that package common data transforms and models. While these reduce engineering for their domains, they do not generalize readily across methods or guarantee environment capture and data versioning by default.
Libraries such as MLXP [arbel_mlxp_2024] focus on lightweight, code-centric management of configurations, runs, and results to reduce overhead and encourage replicable practice. However typically rely on external mechanisms for full environment capture or dataset and and provenance management versioning. Considering this, exact replay remains a user-assembled composition of tools.
@@ III Problem statement
This work aims to design and build a reproducible, low-code experimentation suite for radio localization without precluding advanced experimentation. Today, no widely used approach makes reproducibility the default while also keeping implementation effort low and preserving a clear path to customization. In Section I, we set three goals: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) easy customization. Each goal carries concrete implications for how the system should behave.
@@ IV Proposed framework
To address the problem stated in Section III, we propose a low-code framework LOCALIZE for radio localization that makes reproducible experimentation the default while preserving a clear path to advanced use. The design maps to the requirements in Section III: R1: low-code operation, R2: reproducibility by default, R3: a common evaluation layer, R4: built-in extensibility, and R5: fast setup and iteration.
@@ V-B Language and evnironment
We considered Python, R, Java, and MATLAB for the core runtime. Although MATLAB still appears in localization work [matlab_survey], recent studies and toolchains lean heavily on Python [Roy2021MLSurvey, python_survey]. We chose Python for its mature ML ecosystem (NumPy/Pandas, scikit-learn, TensorFlow/Keras) and smaller coding footprint. The framework is distributed as a standard Python repository. Users clone it and create the environment from an environment.yaml file. Environments are managed with Conda (or Mamba), and dependencies are pinned in the environment file so installations are consistent across machines.
@@ V-E Preconfigured datasets
The repository includes 5 preconfigured datasets and their corresponding dataset-specific stages (Prepare, Featurize, and Split with several different cross-validation strategies) and experiment templates: Lumos5G [narayanan_lumos5g_2020] and UMU datasets for cellular based localization [asensio-garriga_university_2025], LOG-a-TEC dataset for Bluetooth low-energy based localization [morano_outdoor_2022], CTW 2019 [arnold_novel_2019] and CTW 2020 [gauger_massive_2020] datasets for Wifi based localization . For experimenting with any of these datasets, a user can copy a template, edit params.yaml, and run a study without writing glue code or configuration boilerplate.
@@
Feature/Platform LOCALIZE WEKA [weka_book] Orange3 [orange_pdf] Jupyter notebook [granger_jupyter_2021] Ludwig [ludwig_pdf]
@@ VI-A Cross–platform qualitative comparison
As the first step in the evaluation of LOCALIZE, we identify the closest platforms that are strong candidates to enable similar functionality as the proposed framework and dive deeper to understand specific differences. We identify WEKA [weka_book], Orange3 [orange_pdf], Ludwig [ludwig_pdf], and Jupyter notebooks in JupyterLab [granger_jupyter_2021] as the closest relevant platforms.
@@ VI-B Materials
All the quantitative experiments were carried out partly on the UMU cellular localization [morano_outdoor_2022] dataset and primarily on the LOG-a-TEC Bluetooh localization dataset [asensio-garriga_university_2025].
@@ VI-C Lines of code study
For each change, we counted lines of code (LOC) added and LOC removed separately (two bars per platform), excluding blank lines, comments, and lines containing only structural tokens (e.g., brackets/braces/commas).c
@@ VII-C Scalability on LOG-a-TEC
VIII Limitations and Future Work
While LOCALIZE successfully meets its design goals, certain limitations present clear opportunities for future development. The framework is primarily designed and optimized for tabular data. Although its configuration system can reference other data types, such as the tensors in the CTW 2019 dataset, the internal pipeline components are not guaranteed to be compatible. Future work should focus on implementing modular, modality-specific data handlers to natively support a wider range of inputs. Furthermore, the system’s automation is currently limited to hyperparameter optimization. While AutoKeras can tune a model’s hyperparameters, the neural network architecture itself remains static. Integrating Neural Architecture Search (NAS) techniques would be a significant extension, enabling the automatic discovery of novel model topologies.
From a software architecture perspective, the implementation is tightly coupled to its current machine learning backends, specifically scikit-learn with GridSearchCV and TensorFlow via AutoKeras. This design makes it challenging to incorporate alternative libraries, such as PyTorch, or more flexible optimization frameworks such as Optuna and Ray Tune, which were considered during development. A key direction for future work would involve refactoring the training and evaluation logic into more granular, interchangeable components. This would allow users to seamlessly plug in different backends and optimizers, greatly expanding the framework’s flexibility and long-term utility.

### v3

# A Configuration-First Framework for Reproducible, Low-Code Machine Learning: a Localization Use Case

@@ Abstract
As machine learning (ML) becomes the backbone of critical applications—ranging from logistics to healthcare it becomes even more important to keep the results credible and comparable. Everyday workflows should make rigorous experiment specification and exact repeatability the default, without blocking advanced experimentation. However, in practice, researchers face a three-way gap that could be filled by a framework that offers (i) low coding effort for end-to-end studies, (ii) reproducibility by default including versioned code/data/configurations, controlled randomness, isolated runs, recorded artifacts, and (iii) built-in extensibility so new models, metrics, and stages can be added with minimal integration effort. Existing tools rarely deliver all three for machine learning in general and localization workflows, supporting location-based services, in particular. In this paper we introduce a low-code, configuration-first framework in which 1) experiments are declared in human-readable configuration, 2) a workflow orchestrator runs standardized pipelines from data preparation to reporting, and 3) all artifacts, such as datasets, models, metrics, and reports are versioned. We instantiate the framework as LOCALIZE with pre-configured, versioned datasets that reduce initial setup and boilerplate speeding up model development and evaluation. The design, with clear extension points, lets experts add components without reworking the infrastructure. In a qualitative comparison and a head-to-head study against a plain Jupyter notebook baseline, we show that the framework reduces authoring effort while maintaining comparable runtime and memory behavior. Furthermore, using a Bluetooth low-energy dataset we show that scaling across training data ($1\times$ to $10\times$) keeps orchestration overheads bounded as data grows. Overall, the framework, validated for a localization use case, makes reproducible machine-learning-based model development practical, accessible, and extensible.
@@ I Introduction
Background and motivation. The engineering of modern data-driven systems has undergone a paradigm shift. As machine learning (ML), comprised of classical as well as deep learning methods, becomes the backbone of critical applications—ranging from logistics to healthcare—the software engineering (SE) community faces a growing challenge: the reproducibility crisis in research software [19, 11]. While ML provides powerful mappings from raw data to actionable insights, the development process remains largely ad-hoc, characterized by fragile codebases where small changes in global state or environment lead to inconsistent results [5]. This issue is particularly acute in radio-based localization powering the majority of location-based applications such as navigation, logistics, social applications, entertainment, and healthcare [13]. As the radio-based localization research and engineering is transitioning from rigid analytical models to ML powered ones [23], the research and development tools and pipelines that must evolve to support this evolution.
@@
Gaps in current practice. Moving from analytical to machine learning approaches has shown promise in many settings, but raises practical challenges for everyday research. The complexity and heterogeneity of ML pipelines raise the barrier to entry as shown in a recent exploration of AI system development [7]. The study highlights that the stochastic nature of ML introduces a unique cognitive load on developers that traditional SE practices do not adequately address. GUI/low-code platforms reduce coding effort, but can constrain advanced experimentation and fine-grained control [25, 6]. Conversely, code-centric workflows provide maximum flexibility but, absent dedicated tooling, place the burden of reproducibility, environment control, and consistent evaluation on the researcher [19, 11, 12, 5, 10]. In many research studies, including localization [23], the primary novelty is a new model or training method evaluated on public datasets. Onboarding data and recreating baselines becomes setup overhead rather than the research focus. The broader ML literature shows that small changes to preprocessing, hyperparameters, software versions, or random seeds can alter conclusions and hinder fair comparison [19, 11]. Beyond statistical variance, exact repeatability is fragile: hidden global state, non-deterministic operations, data-order effects, and environment or hardware differences can change outputs even when seeds and settings are fixed [5, 12, 18, 27]. Such issues are easy to miss and hard to diagnose.
@@
This challenge represents a critical bottleneck in many scientific domains now adopting machine learning [11]. Fields such as computational biology and materials science face a similar gap where the transition from traditional models to complex, data-driven pipelines strains existing research practices [11]. The lack of standardized, reproducible, and extensible workflows leads to duplicated effort, fragile results, and a high barrier to entry for domain experts who are not also software engineers [24]. Therefore, the framework presented in this paper addresses a fundamental software engineering problem: how to develop research-specific experimentation environments that enforce scientific rigor by default without sacrificing the agility needed for rapid experimentation.
@@
Proposed approach and contributions. This paper addresses the software engineering problem of how to design ML experimentation environments that enable: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) flexibility that is built in, so advanced experimentation can be conducted without fighting the tool-chain. We adopt a configuration-first approach in which experimental intent is explicitly declared, execution is orchestrated in isolated stages, and all artifacts are systematically tracked. This design follows the principle of engineering within boundaries [17] for non-deterministic systems, treating reproducibility not as an optional feature but as an architectural invariant. We instantiate this approach in LOCALIZE [^1] https://github.com/sensorlab/localize, a low-code, configuration-first framework that makes ML-based experimentation practical, accessible, and reproducible by reducing initial effort, enabling novice and expert use, and supporting expert extensibility while promoting sound experimental practices. While the reference implementation targets radio localization workflows, the architectural principles and empirical findings are intended to inform the design of reproducible ML research software more broadly. The contributions of this paper presents are:
@@
  • We introduce a low-code, configuration-first framework that bridges ease of use with scientific rigor by making reproducibility its default operating mode, integrating version control, execution isolation, and transparent artifact tracking.
@@
  • We design the framework to enforce consistent, comparable evaluation by applying a standardized set of metrics and reporting procedures across all methods and datasets, eliminating glue code and improving the credibility of results.
@@
Collectively, these contributions position the work not only as a domain-specific tool, but as a case study in the engineering of reproducible ML research environments, with design principles and empirical insights applicable to other data-driven scientific domains. In the remainder of this paper, we review related work in Section II, outline the problem and requirements in Section III, describe the proposed framework in Section IV, present a reference implementation in Section V, and evaluate it against established platforms in Section VII. Finally, Section IX concludes the paper.
@@ II-A The Challenge of Reproducibility in Machine Learning
A consistent lesson from the literature is that ML results can be sensitive to seemingly minor choices, so reproducibility requires attention to both variability and exact repeatability. Bouthillier et al. [5] quantify sources of variability in benchmarks and show that initialization, data sampling and order, and hyperparameter settings all affect outcomes. They recommend practices such as repeated runs and clearer reporting to separate algorithmic gains from noise. Building on this empirical picture, Henderson et al. [12] demonstrate that non-determinism in benchmark environments and implementation details can materially change results, arguing for stricter methodology and documentation. Furthermore, [27] recently formalized the concept of ’reproducibility smells’ defined as patterns in the development lifecycle, such as hidden global states and lack of automated artifact tracking, that lead to fragile and non-deterministic results. Collectively, these studies establish that achieving computational reproducibility requires meticulous control over the entire experimental pipeline.
In response, subsequent work outlines reporting and release requirements. Heil et al. [11] introduce tiered levels of computational reproducibility (bronze/silver/gold) with concrete requirements covering code, data, captured environments, and automation. At the community level, Pineau et al. [19] describe the NeurIPS reproducibility program, checklists, a code policy, and a community challenge, which raised expectations for transparency across submissions.
Despite these advances, practical support in general ML platforms remains uneven. Gundersen et al. [10] report that reproducing published results often requires non-trivial user effort. This suggests a clear need for research workflows that embed controls, documentation, and repeatable execution as default, integral features rather than optional, post-hoc additions.
@@ II-B General-Purpose Experimentation Platforms
Several general-purpose platforms aim to simplify ML experimentation while only partly alleviating the cognitive load of AI development [7]. They typically fail to resolve the tension between usability, reproducibility, and extensibility, forcing researchers to make compromises. For example, graphical tools like WEKA [25] and visual programming environments like Orange3 [6] excel at lowering the barrier to entry in machine learning research in different domains. They allow for rapid prototyping without extensive coding. However, they externalize the burden of reproducibility, as rigorously capturing an end-to-end experiment, including precise library versions and sources of non-determinism, requires manual effort outside the core workflow.
Declarative, code-centric platforms like Ludwig [15] advance this goal by using configuration files to define deep learning experiments. This approach streamlines workflows and supports extensibility. While powerful for neural networks, it is not primarily designed for the classical machine learning pipelines common in localization studies and, similar to graphical tools, leaves environment and data versioning as an external concern.
These platforms assist with parts of the problem, namely addressing low-code use or standardizing certain workflows, but they do not jointly provide a solution that is reproducible by default, accessible via a low-code interface, and flexible enough for advanced experimentation.
@@ II-C Other Related Efforts
Specialized tools also contribute valuable insights, typically focusing on specific domains or aspects of experimental management. Some efforts focus on automating workflows for a specific domain, such as CSI fingerprinting localization [22] or multivariate time-series forecasting [26], that package common data transforms and models. While these reduce engineering for their domains, they do not generalize readily across methods or guarantee environment capture and data versioning by default.
Libraries such as MLXP [1] focus on lightweight, code-centric management of configurations, runs, and results to reduce overhead and encourage replicable practice. However typically rely on external mechanisms for full environment capture or dataset and and provenance management versioning. Considering this, exact replay remains a user-assembled composition of tools.
@@ III Problem statement
This work aims to design and build a reproducible, low-code experimentation suite for advanced ML experimentation. Today, no widely used approach makes reproducibility the default while also keeping implementation effort low and preserving a clear path to customization. In Section I, we set three goals: (i) reproducibility by default, (ii) ease of use (low-code with minimal boilerplate), and (iii) easy customization. Each goal carries concrete implications for how the system should behave.
@@ IV Proposed framework
To address the problem stated in Section III, we propose LOCALIZE, a low-code framework that makes reproducible experimentation the default while preserving a clear path to advanced use. The design maps to the requirements in Section III: R1: low-code operation, R2: reproducibility by default, R3: a common evaluation layer, R4: built-in extensibility, and R5: fast setup and iteration.
@@ V-B Language and evnironment
We considered Python, R, Java, and MATLAB for the core runtime. Although MATLAB still appears in localization work [14], recent studies and toolchains lean heavily on Python [21, 20]. We chose Python for its mature ML ecosystem (NumPy/Pandas, scikit-learn, TensorFlow/Keras) and smaller coding footprint. The framework is distributed as a standard Python repository. Users clone it and create the environment from an environment.yaml file. Environments are managed with Conda (or Mamba), and dependencies are pinned in the environment file so installations are consistent across machines.
@@ V-E Preconfigured datasets
The repository includes 5 preconfigured datasets and their corresponding dataset-specific stages (Prepare, Featurize, and Split with several different cross-validation strategies) and experiment templates: Lumos5G [16] and UMU datasets for cellular based localization [3], LOG-a-TEC dataset for Bluetooth low-energy based localization [4], CTW 2019 [2] and CTW 2020 [8] datasets for Wifi based localization . For experimenting with any of these datasets, a user can copy a template, edit params.yaml, and run a study without writing glue code or configuration boilerplate.
@@
Feature/Platform LOCALIZE WEKA [25] Orange3 [6] Jupyter notebook [9] Ludwig [15]
@@ VI-A Cross–platform qualitative comparison
As the first step in the evaluation of LOCALIZE, we identify the closest platforms that are strong candidates to enable similar functionality as the proposed framework and dive deeper to understand specific differences. We identify WEKA [25], Orange3 [6], Ludwig [15], and Jupyter notebooks in JupyterLab [9] as the closest relevant platforms.
@@ VI-B Materials
All the quantitative experiments were carried out partly on the UMU cellular localization [4] dataset and primarily on the LOG-a-TEC Bluetooh localization dataset [3].
@@ VI-C Lines of code study
For each change, we counted lines of code (LOC) added and LOC removed separately (two bars per platform), excluding blank lines, comments, and lines containing only structural tokens (e.g., brackets/braces/commas).
@@ VII-C Scalability on LOG-a-TEC
VIII Threats to Validity
We discuss threats to validity following established empirical software engineering guidelines, reflecting limitations identified during the design and evaluation of the LOCALIZE framework.
Internal validity concerns whether observed effects are attributable to the framework rather than confounding factors. The quantitative comparison against a Jupyter notebook baseline relies on functionally equivalent pipelines implemented by the authors. Although both implementations used identical algorithms, datasets, hyperparameters, and evaluation procedures, some implementation bias may remain. To mitigate this risk, both pipelines were validated to produce identical outputs under identical configurations, and execution isolation was enforced in the notebook baseline by running each experiment in a fresh process. Runtime and memory measurements were conducted in a controlled environment with consistent hardware and software configurations, and results were averaged across repeated runs. Nevertheless, transient system effects and library-level nondeterminism may still introduce minor variability, which is inherent to ML experimentation and was minimized but not entirely eliminated.
Construct validity addresses whether the selected metrics capture the intended constructs. Developer effort was approximated using lines of code (LOC) added and removed. While LOC is objective and reproducible, it does not capture cognitive effort, debugging time, learning curve, or error proneness. We therefore interpret LOC differences as indicators of relative engineering overhead rather than definitive productivity measures. This limitation is partially mitigated by focusing on controlled, incremental experimental changes representative of common research tasks. The qualitative cross-platform comparison relies on ordinal ratings derived from documentation review and hands-on use, which are inherently subjective. To reduce bias, evaluation criteria were explicitly defined, and scores were capped when functionality required non-standard extensions or external tooling.
External validity concerns generalizability beyond the studied context. The reference implementation and empirical evaluation focus on radio localization workflows using tabular datasets derived from wireless measurements. While the underlying architectural principles—configuration-first design, execution isolation, artifact lineage, and evaluation invariance—are domain-agnostic, the concrete pipeline components are optimized for this class of problems. Consequently, the quantitative results may not directly generalize to workflows involving other data modalities (e.g., images, graphs, or streaming data) or to online learning, continuous deployment, or production MLOps scenarios. We mitigate this threat by positioning LOCALIZE as a reproducibility-by-design research framework and a case study in engineering ML experimentation environments, rather than as a universal solution.
Reproducibility validity concerns whether the reported results can be independently replicated. All experiments used fixed random seeds, versioned code, configuration, and data, and isolated execution environments. Nonetheless, exact numerical reproducibility may still be affected by differences in hardware architectures, operating systems, or low-level numerical libraries, particularly for deep learning components. This limitation is intrinsic to current ML software stacks. To mitigate this risk, the complete implementation, configurations, and datasets are publicly available, enabling independent replication and inspection.
Design and scope limitations arise from choices prioritizing simplicity and reproducibility. The framework is optimized for tabular data and classical ML workflows, with neural models supported primarily through AutoKeras. Supporting additional backends (e.g., PyTorch), optimization frameworks (e.g., Optuna or Ray Tune), or data modalities would require further architectural refactoring. These limitations do not undermine the validity of the results but delineate the current scope of applicability.